idme7:

Grant Valley School has hired a fleet of buses and vans to transport 96 students to a picnic. Each bus can seat 31 people, and each van can seat 9 people. The school authorities want to have three teachers traveling in each bus and one teacher in each van. It will take to accommodate 107 people, including the students and teachers, with no seats left vacant.

jhonyy9:

what is the question here ?

Extrinix:

@jhonyy9 wrote:
what is the question here ?
It’s asking how many buses and vans will it take to accommodate 107 people, leaving no spaces open

Extrinix:

I’m thinking something like this? \(\sf{107 = 31(x) + 9(y)}\) Where \(\sf{x}\) equals the amount of buses and \(\sf{y}\) equals the amount of vans. Am I looking at this correct, @jhonyy9 ?

jhonyy9:

yes but there is a second equation too

jhonyy9:

,, The school authorities want to have three teachers traveling in each bus and one teacher in each van."

Extrinix:

Ah so \(\sf{107 = 31(3 + x) + 9(1 + y)}\)?

Extrinix:

No because that would be different, I need to swap out the variable placements \(\sf{96 = x(31-3)+y(9-1)}\) That’s what I’m thinking… ?

jhonyy9:

@extrinix wrote:
Ah so \(\sf{107 = 31(3 + x) + 9(1 + y)}\)?
I think this 31 will be 28+3 and this 9 will be 8+1

Extrinix:

@jhonyy9 wrote:
@extrinix wrote:
Ah so \(\sf{107 = 31(3 + x) + 9(1 + y)}\)?
I think this 31 will be 28+3 and this 9 will be 8+1
I corrected myself because 107 is the total amount (students and teachers), so we could just remove the amount of teachers from the equation and from the total slots on each vehicle type

Extrinix:

That’s where this equation came from \(\sf{\downarrow}\) \(\sf{\downarrow}\)

@extrinix wrote:
No because that would be different, I need to swap out the variable placements \(\sf{96 = x(31-3)+y(9-1)}\) That’s what I’m thinking… ?

jhonyy9:

this 31 = 28 students + 3 teachers in a bus and 9 = 8 students +1 teacher in a van

surjithayer:

96+3x+y=107 3x+y=107-96=11 31x+9y=107 31x+9(11-3x)=107 31x+99-27x=107 4x=107-99=8 x=8/4=2 2(3)+y=11 y=11-6=5 number of buses=2 number of vans=5

surjithayer:

here number of buses=x number of vans=y